# A Generative Framework for Predictive Modeling of Multiple Chronic Conditions Using Graph Variational Autoencoder and Bandit-Optimized Graph Neural Network

### v1

@@ 4 Discussion
For 15 iterations, the $\varepsilon$-Greedy algorithm displays a median accuracy of approximately 0.72 with a moderate interquartile range (IQR), indicating consistent performance with some variability. The Multi-Armed Bandit (MAB) algorithm shows the lowest median accuracy, around 0.6, with a smaller IQR, reflecting consistently lower performance but less variability. Conversely, the Contextual Bandit (CB) algorithm achieves a median accuracy close to 0.73 with a larger IQR, suggesting better performance despite greater variability.
At 20 iterations, $\varepsilon$-Greedy maintains a median accuracy of around 0.72 but exhibits a slight increase in variability. MAB continues to have the lowest median accuracy, around 0.6, with an increased IQR, indicating more variability. CB again shows the highest median accuracy, approximately 0.83, with an even larger IQR, reflecting its continued superiority despite higher variability.
With 25 iterations, $\varepsilon$-Greedy shows a slight improvement in median accuracy and reduced variability, suggesting better consistency. MAB’s performance remains similar to previous iterations, with low median accuracy and a small IQR. CB achieves the highest median accuracy, around 0.83, with a reduced IQR compared to previous iterations, indicating improved consistency.

### v2

@@ 4 Discussion
For 15 iterations, the $\varepsilon$-Greedy algorithm displays a median accuracy of approximately 0.72 with a moderate interquartile range (IQR), indicating consistent performance with some variability. The Multi-Armed Bandit (MAB) algorithm shows the lowest median accuracy, around 0.6, with a smaller IQR, reflecting consistently lower performance but less variability. Conversely, the Contextual Bandit (CB) algorithm achieves a median accuracy close to 0.73 with a larger IQR, suggesting better performance despite greater variability. At 20 iterations, $\varepsilon$-Greedy maintains a median accuracy of around 0.72 but exhibits a slight increase in variability. MAB continues to have the lowest median accuracy, around 0.6, with an increased IQR, indicating more variability. CB again shows the highest median accuracy, approximately 0.83, with an even larger IQR, reflecting its continued superiority despite higher variability. With 25 iterations, $\varepsilon$-Greedy shows a slight improvement in median accuracy and reduced variability, suggesting better consistency. MAB’s performance remains similar to previous iterations, with low median accuracy and a small IQR. CB achieves the highest median accuracy, around 0.83, with a reduced IQR compared to previous iterations, indicating improved consistency.
@@
4.1 Model Generalizability
To evaluate the generalizability of our framework beyond the original study population, we performed an external validation using a more generalized synthetic dataset. This dataset was evaluated using the best-performing combination with the CB algorithm as the reinforcement learning model. Given the lack of publicly available datasets with a structure and feature set comparable to the CCHC, we created a synthetic dataset that not only preserved key variable relationships but also incorporated additional risk factors. These new features were designed to enhance model robustness by reflecting a broader range of health conditions and patient characteristics.
To ensure the synthetic dataset integrity, we conducted a distribution analysis using the distfit[54] Python library on key variables such as age (Dagum distribution), systolic and diastolic blood pressure (Johnson’s SU-distribution), gender (Cauchy distribution), among others. This process allowed us to generate realistic risk factor distributions while maintaining statistical consistency with the original CCHC dataset. The synthetic dataset was constructed in a way that retained essential dependencies while introducing controlled variability, ensuring a more rigorous test of the model’s ability to generalize.
The models were then trained and evaluated on this synthetic dataset, with predictive performance measured using precision, recall, F1-score, and AUC-ROC, see Table 5. These metrics provide a comprehensive assessment of the model’s capability to handle variability in healthcare data. The results of this validation confirm the model’s ability to generalize beyond the initial cohort, reinforcing the reliability of our approach for predictive modeling of MCC in diverse populations.
Table 5: Validating proposed model with synthetic dataset
GVAE-GNN 15 Iterations 20 Iterations 25 Iterations
Avg. Test Accuracy 0.7809 0.7965 0.8165
F1-Score 0.7813 0.7945 0.8037
Recall 0.7818 0.7941 0.8035
Precision 0.7865 0.8009 0.8023
AUC 0.9823 0.9839 0.9887